# Prompto: An open source library for asynchronous querying of LLM endpoints

### v1

@@ 3.4 Adding new APIs and models to prompto
4 Experiments
We provide some illustrative examples [^23] All experiments were ran on a 2021 MacBook Pro with M1 Pro and 32 GB of memory. of using our prompto library and compare against a traditional synchronous approach to querying LLM endpoints. In each of the following examples, we consider the task of generating instruction-following data using the Self-Instruct approach of Wang et al. (2022) and Taori et al. (2023). We take a sample of $100100$ prompts from the instruction-following data [^24] https://github.com/tatsu-lab/stanford_alpaca/blob/main/alpaca_data.json from Taori et al. (2023) and apply the same prompt template. We then use these as prompt inputs to different models using prompto. Full details of how we obtained a sample of prompts along with notebooks to run each of the following experiments can be found at https://alan-turing-institute.github.io/prompto/examples/system-demo/.
4.1 Querying LLM endpoints asynchronously vs synchronously
@@
4.2 Querying different LLM endpoints in parallel vs synchronously
As mentioned in Section 3.2, prompto supports sending requests to different APIs in parallel for greater efficiency gains. Further, the user can easily specify how many queries to send to each API or model per minute. To illustrate, we consider the same APIs and models as in Section 4.1, but query the models in parallel. We compare against the baseline synchronous approach which we expect to be close to the sum of the individual run-times of the synchronous approaches in Table 1. We report run-times of the two approaches to obtain a total of $300300$ prompts ($100100$ to each model/API) in Table 2.
We observe a $22$ times speedup when using prompto with parallel processing. The prompto run-time is close to how long it took to process the Ollama requests in Section 4.1 which is expected since the run-time of querying different APIs or models in parallel is simply dictated by the slowest API or model to query.
@@ 4.2 Querying different LLM endpoints in parallel vs synchronously
4.3 Querying different models from the same endpoint in parallel vs synchronously
@@
For each model and with parallel processing of the models, using prompto offers a signficant speedup in obtaining responses. Focusing on the synchronous run-times, we can see GPT-4o and GPT-4 are slower to obtain responses for than GPT-3.5-Turbo. Comparing using prompto with parallel processing against synchronously querying each model, we obtain an approximately $3535$ times speedup with prompto. As with the previous experiment in Section 4.2, the run-time of prompto with parallel processing is roughly the time to query the slowest model, GPT-4 in this case.
@@
5 Conclusions and Future Work
We present prompto, an open source library designed to facilitate researchers to efficiently query large language models which reside on proprietary or self-hosted API endpoints. We test our pipeline by comparing against a traditional synchronous approach to querying LLM endpoints and show significant speedups, especially when querying different models either on separate LLM endpoints (Section 4.2) or available on the same API service (Section 4.3). Using prompto is simple and all inputs/prompts can be easily defined in a single JSONL file, which can encompass queries from various APIs and models. Outputs are then stored in output JSONL files (i.e. an experiment file where each prompt dictionary has a completed "response" key for the model output) allowing researchers to easily share experiment outputs promoting reproducibility. Consequently, our library enables faster experimentation and evaluation.
Specifically regarding evaluation applications, we hope in future versions to improve upon functionality for automatic evaluation of responses. Currently, there are commands to generate new input experiment JSONL files for automatic model graded evaluations like the LLM-as-a-judge approach of Zheng et al. (2024) given an output file. However, there may be improvements on this to allow further custom evaluations.
prompto is an ongoing open source project and is set up such that adding new APIs is simple. We welcome contributions from the community to ensure support for a wide range of LLM endpoints and new features. In particular, while prompto currently supports constructing prompts and inputs for multimodal LLMs for some APIs such as the Gemini and VertexAI. Extending multimodal support for other APIs and improving this process is a potential area for improvement to foster their usage and comparisons in experimental research settings.
Acknowledgments
We thank Liam Burke-Moore, Yi-Ling Chung, Florence Enock, Kobi Hackenburg & Tvesha Sippy for useful discussions on this library. This work was partially supported by the Ecosystem Leadership Award under the EPSRC Grant EPX03870X1, The AI Safety Institute & The Alan Turing Institute.

### v2

@@ 3.6 Adding new APIs and models to prompto
4 Conclusions and Future Work
We present prompto, an open source library designed to facilitate researchers to efficiently query large language models which reside on proprietary or self-hosted API endpoints. Using prompto is simple and all inputs/prompts can be easily defined in a single JSONL file, which can encompass prompts for various APIs and models. Outputs are then stored in output JSONL files allowing researchers to easily share experiment outputs promoting reproducibility. Consequently, our library enables faster experimentation and evaluation. prompto is an ongoing open source project and we welcome contributions from the community to ensure support for a wide range of LLM endpoints and new features.
There are a number of interesting avenues for extending prompto. In the future, we hope to continue focusing on extending the evaluation pipelines discussed in Section 3.4. For evaluation, integration of larger-scale NLP pipelines can be useful in order to efficiently apply more advanced scoring functions such as toxicity classifiers and other safety risk classifiers for automatic safety evaluation. Furthermore, integrating prompto with data visualisation libraries can bring additional plotting features to the library to enable deeper exploration of model performance.
Another direction of work is to extend the rephrasing pipelines in prompto (see Section 3.5). A particular area of interest is extending the pipeline for translation of prompts. This can be useful for synthetic data generation applications. The current pipeline enables the use of LLMs for translation which have demonstrated remarkable potential in handling multilingual machine translation (MMT) (Zhu et al., 2024), however extending this pipeline to dedicated MMT models, such as NLLB (Costa-jussà et al., 2022; NLLB Team, 2024), could particularly be effective for translating to lower resourced languages. Moreover, a rephrasing pipeline can also be important in evaluation settings to generate a wider set of prompts to test a model. However, further research is required in order to do this effectively.
Lastly, future work includes expanding our tool’s accessibility by developing direct local model integration capabilities. The current implementation requires setting up an API endpoint for querying, but we envision simplifying this process by enabling direct model instantiation (e.g., using transformers (Wolf et al., 2020)). While toolkits for deploying and serving LLMs, such as text-generation-inference or vLLM (Kwon et al., 2023), already simplify setting up efficient endpoints, this approach would reduce overhead for researchers who wish to query local models and utilise offline batched inference, allowing for more streamlined experimentation and rapid prototyping.
Acknowledgments
We thank Yi-Ling Chung, Florence Enock, Kobi Hackenburg for useful discussions on this library. This work was partially supported by the Ecosystem Leadership Award under the EPSRC Grant EPX03870X1, The AI Safety Institute & The Alan Turing Institute.
Appendix A Experiments
We provide some illustrative examples [^21] All experiments were ran on a 2021 MacBook Pro with M1 Pro and 32 GB of memory. of using our prompto library and compare against a traditional synchronous approach to querying LLM endpoints. We show that our our pipeline provides significant speedups, especially when querying different models either on separate LLM endpoints (Appendix A.2) or available on the same API service (Appendix A.3).
In each of the following examples, we consider the task of generating instruction-following data using the Self-Instruct approach of Wang et al. (2022) and Taori et al. (2023). We take a sample of $100100$ prompts from the instruction-following data [^22] https://github.com/tatsu-lab/stanford_alpaca/blob/main/alpaca_data.json from Taori et al. (2023) and apply the same prompt template. We then use these as prompt inputs to different models using prompto. Full details of how we obtained a sample of prompts along with notebooks to run each of the following experiments can be found at https://alan-turing-institute.github.io/prompto/examples/system-demo/.
Further examples including examples with prompting multimodal models with images and videos can be found on our GitHub repo and documentation\@footnotemark.
A.1 Querying LLM endpoints asynchronously vs synchronously
@@
A.2 Querying different LLM endpoints in parallel vs synchronously
As mentioned in Section LABEL:subsec:experiment_run, prompto supports sending requests to different APIs in parallel for greater efficiency gains. Further, the user can easily specify how many queries to send to each API or model per minute. To illustrate, we consider the same APIs and models as in Section A.1, but query the models in parallel. We compare against the baseline synchronous approach which we expect to be close to the sum of the individual run-times of the synchronous approaches in Table 1. We report run-times of the two approaches to obtain a total of $300300$ prompts ($100100$ to each model/API) in Table 2.
We observe a $22$ times speedup when using prompto with parallel processing. The prompto run-time is close to how long it took to process the Ollama requests in Section A.1 which is expected since the run-time of querying different APIs or models in parallel is simply dictated by the slowest API or model to query.
@@ A.2 Querying different LLM endpoints in parallel vs synchronously
A.3 Querying different models from the same endpoint in parallel vs synchronously
@@
For each model and with parallel processing of the models, using prompto offers a signficant speedup in obtaining responses. Focusing on the synchronous run-times, we can see GPT-4o and GPT-4 are slower to obtain responses for than GPT-3.5-Turbo. Comparing using prompto with parallel processing against synchronously querying each model, we obtain an approximately $3535$ times speedup with prompto. As with the previous experiment in Section A.2, the run-time of prompto with parallel processing is roughly the time to query the slowest model, GPT-4 in this case.